OpenStudy (anonymous):

First derivative of (1+x^2) Arctan x-x

OpenStudy (primeralph):

What?

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}(1+x^2)\arctan x-x\]

OpenStudy (primeralph):

What's being operated upon?

OpenStudy (anonymous):

inverse trigo

OpenStudy (anonymous):

the answer was 2x arctan x.. and still, no idea deriving the right answer

OpenStudy (primeralph):

We need the right question to give the right answer.

hartnn (hartnn):

\(\Large\bf\sf \frac{d}{dx}(1+x^2)\arctan x-x\) does give 2x arctan x

hartnn (hartnn):

did u try product rule for first term

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

ah! yes

hartnn (hartnn):

heard of product rule ? \(\huge (uv)' = u'v+uv'\) here u= 1+x^2 v = arctan x there you go :)

OpenStudy (anonymous):

i got 2x arctan x-x.. and.. it wasnt x-x,

hartnn (hartnn):

how come you get the 2nd term as -x ?? shouldn't it be -1 ?

OpenStudy (primeralph):

You guys should really be more clear about the derivatives.

hartnn (hartnn):

i mean +1

OpenStudy (anonymous):

what i did first was (1+x^2) (x-x dy/dx / 1-2x^2) + (2x) (arc tan x-x)

OpenStudy (primeralph):